# Stabilizer Entanglement as a Magic Highway

### v2

@@ A.2 A useful quantity and its approximations
where $P_{A}$ is a Pauli string supported on subsystem $A$. In the above definition, we also abuse notation by using $\mathbb{S}_{A}$ to represent an equal weight summation of all elements in $\mathbb{S}_{A}$: $\sum_{g\in\mathbb{S}_{A}}g$. The connection between the above quantity and $\overline{Y}$ will be explained in detail in the next subsection. Basically it naturally arises when computing ${\rm Tr}(PU_{A}\rho U_{A}^{\dagger})^{4}$, with $\rho$ represented as a projector onto the stabilizer subspace.
Eq. (10) can be represented using the diagrammatic method introduced in Sec. A.1 as follows:
Figure 5: (a) Leading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$. There are $576{4!}^{2}=$ 576 diagrams in total. If $P_{A}\neq I_{A}$, then $\text{Tr}(P_{A})=\text{Tr}(P_{A}^{3})=0$, and there are 216 diagrams with non-zero contributions. (b) Subleading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$.
@@
A.3 Derivation of the key results presented in the main text
In this subsection, we provide a detailed derivation of Eqs. (1), (4) and (5) in the main text.
A.3.1 Eq. (1)
We start with Eq. (1). The key to our derivation is to represent the initial state as a projector onto the stabilizer subspace, and a decomposition of the stabilizer group of the full sytem in terms of the cosets of stabilizer subgroups of subsystems $A$ and $B$. Let $\mathbb{S}_{AB}$ denote the stabilizer group of the initial state. Then initial density matrix can be written as a projector
@@ A.3 Derivation of the key results presented in the main text
where $a_{i}\notin\mathbb{S}_{A}$ and $b_{i}\notin\mathbb{S}_{B}$ are the logical operators of subsystems $A$ and $B$, respectively. To ensure the proper commutation relation, the logical operators must be paired up into $(a_{i},b_{i})$ such that: (i) if $[a_{i},a_{j}]=0$, then $[b_{i},b_{j}]=0$; or (ii) if $\{a_{i},a_{j}\}=0$, then $\{b_{i},b_{j}\}=0$. While it is not hard to convince oneself that the above decomposition is correct for pure states with ${\rm dim}(\mathbb{S}_{AB})=N$, let us check that the dimensions work out. First of all, since the entanglement between $A$ and $B$ is $E$, there are precisely $E$ logical qubits encoded in both subsystems, with $4^{E}$ logical operators in each subsytem, and hence $4^{E}$ terms in the decomposition. Secondly, we can count the total number of group elements in $\mathbb{S}_{AB}$ from the right-hand side of the decomposition. We have $4^{E}\times 2^{{\rm dim}(\mathbb{S}_{A})}\times 2^{{\rm dim}(\mathbb{S}_{B})}=2^{|A|+|B|}=2^{N}$, as it should. We will give a concrete example of such a decomposition in Appendix B.
@@
Armed with the above representation of $\rho_{0}$, we can cast the average value $\overline{Y}$ as:
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{A}}{\mathbb{S}_{AB}}{{U}_{A}}^{\dagger})}}$ (17)
@@
Consider first a fixed Pauli string $P$ and $k$, which corresponds to one term of the summation in $\overline{Y}$. We have the following result:
@@
$=$ $\delta(P_{B},b_{k}\mathbb{S}_{B})\times\begin{cases}3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1-2\cdot{{2}^{E-{|A|}}}+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k=0\\ 3\cdot{{2}^{2{|A|}+4{|B|}-2E}}\left[1+O({{2}^{-E-{|A|}}})\right],&\quad{{P}_{A}}\neq{{I}_{A}},\quad k\neq 0\\ {{2}^{4{|A|}+4{|B|}}},&\quad{{P}_{A}}={{I}_{A}},\quad k=0\\ 0,&\quad{{P}_{A}}={{I}_{A}},\quad k\neq 0\\ \end{cases}$
@@
A.3.2 Eq. (4)
We provide detailed calculations for Eq. (4) in the main text. Since the unitary acting on $\rho_{0}$ takes a factorized form $U=U_{A}\otimes U_{B}$, the average $\overline{Y}$ similarly factorizes into
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}},$ (20)
@@
We now analyze the order of magnitude of each term. Since $\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}\cdot(...)\leq\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}$, where ()(...)( … ) represents any product of $\delta_{1},\delta_{2}$, or $\delta_{4}$ (as additional $\delta$-functions further reduce the number of unconstrained summations over $k_{1},\ldots,k_{4}$), we have
@@
After combining all these contributions and including the multiplicity factors coming from the number of Pauli strings in each case, we obtain the final expression in Eq. (4) of the main text.
A.3.3 Eq. (5)
We now present the calculation details of Eq. (5) in the main text. Let us first recall the setup we use for this case. We consider a tripartite system $A$, $B$, and $C$. Since any tripartite stabilizer state can be transformed into a set of single-qubit states, Bell-pair states, and GHZ states via local Clifford unitaries, we consider initial state consisting of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem. Let us start by explicitly writing down a set of stabilizer generators for this state.
@@
where $l_{AB}^{k}$ and $l_{C}^{k}$ are logical operators on $AB$ and $C$ as before. In this case, since we are dealing with initial state with a factorized structure, it is easy to directly idenfity the logical operators. For example, $\{l_{AB}^{k}\}$ consists of Pauli operators acting on $b_{AC}+b_{BC}$ Bell pairs, as well as Pauli operators acting on $g$ GHZ states, hence $4^{b_{AC}+b_{BC}+g}$ operators in total. As before, to ensure elements across different cosets have the correct commutation relation as stabilizer group elements, we require (i) if $[l_{AB}^{i},l_{AB}^{j}]=0$, then $[l_{C}^{i},l_{C}^{j}]=0$; or (ii) if $\{l_{AB}^{i},l_{AB}^{j}\}=0$, then $\{l_{C}^{i},l_{C}^{j}\}=0$.
The average $\overline{Y}$ can be written as
$\overline{Y}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{AB}\otimes I_{C}}{\mathbb{S}_{ABC}}{U}_{AB}^{\dagger}\otimes I_{C})}}$ (30)
@@
The $\overline{Y}$ is then expressed in Eq. (57):
$2^{5|A|+5|B|+5|C|}\cdot 2^{-4|C|}\cdot 2^{-|C|+b_{AC}+b_{BC}+g}\cdot\overline{Y}$ (57)
@@
Appendix B An example of the decomposition of the stabilizer group
@@
$\quad\quad\mathbb{S}_{A}\otimes\mathbb{S}_{B}\quad\quad$ $XXX\mathbb{S}_{A}\otimes XX\mathbb{S}_{B}$ $IIZ\mathbb{S}_{A}\otimes ZI\mathbb{S}_{B}$ $-XXY\mathbb{S}_{A}\otimes YX\mathbb{S}_{B}$
@@
Appendix C Factorized unitary acting on three subregions
@@
$\overline{Y}=4\cdot 2^{-N}\big{[}1+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{BC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}$ (62)
@@
where once again subleading corrections are of the order $O(2^{-|A|},2^{-|B|},2^{-|C|})$. Similarly to the bipartite result Eq. (4) in the main text, we find that the resulting state also achieves maximal amount of magic, provided that the subsystems are sufficiently entangled. Interestingly, this does require that the $A$, $B$, and $C$ are pairwise entangled. As Eq. (62) suggests, if $A\&B$, $B\&C$ are highly entangled while $A\&C$ are not entangled (i.e. $b_{AB},b_{BC}$ is large, but $b_{AC}$ is zero), the average $\overline{Y}$ still approaches that of a Haar random state.
As a consistency check, we consider a special case of Eq. (62), namely, when only $A\&B$ are entangled, and $C$ is not entangled with $AB$. In this case, we can check that the average $\overline{Y}$ coincides with a product of Eq. (4) in the main text for the bipartite case and a single Haar random state on $C$, which is consistent with our previous results.
@@
$\overline{Y}=\frac{1}{{2^{5N}}}\sum\limits_{P\in P_{N}}\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\alpha_{P_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})\beta_{P_{B}}(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}})\gamma_{P_{C}}(c_{k_{1}},c_{k_{2}},c_{k_{3}},c_{k_{4}}).$ (64)
Now, depending on the operator contents of the Pauli string $P$ on subsystems $A$, $B$ and $C$ (identity or not), we use Eq. (12) and Eq. (13) to evaluate the above expression. We give one example here. Consider when $P_{A}\neq I_{A}$, $P_{B}\neq I_{B}$, and $P_{C}\neq I_{C}$, using Eq. (12) and Eq. (13), we expand Eq. (64) as follows:
@@
Appendix D Magic injection with shallow-depth brickwork circuits
In this section, we replace the Haar random unitary gate acting on $A$ with a shallow-depth brickwork circuit. First, we still consider the initial state shown in Fig. 1(d) of the main text, followed by one layer of local gates acting on subregion $A$. Since the initial state factorizes into single-qubit states and Bell-pair states, the average $\overline{Y}$ also factorizes into a product of contributions from single-qubit and Bell pair states. Thus, it is easy to see that, $\overline{Y}\propto 2^{-E}$ still holds true in this case.
However, if $U_{A}$ is not Haar random, the specific choice of initial state in Fig. 1(d) is no longer the most generic one. To construct the most general stabilizer initial state with tunable entanglement, notice that since any bipartite stabilizer state is local-unitary equivalent to the state in Fig. 1(d), we can simply recover a generic stabilizer state by acting on Fig. 1(d) with random Clifford unitaries $U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}$. Since $U^{\rm Cliff}_{B}$ commutes with subsequent magic injection gates supported on $A$ and leaves $Y$ invariant, we can simply apply $U^{\rm Cliff}_{A}$. We then consider brickwork circuits of varying depths and gate ranges, as summarized in Fig. 6. We then numerically compute the average $\overline{Y}$ corresponding to each circuit architecture, and the results are shown in Fig. 7. We find that the qualitative picture of our analytical result is still valid for the brickwork circuit, in that the average $\overline{Y}$ decays exponentially with $E$. However, notice that for extremely shallow depths (single layer and two layers), the curve bends upward and saturates, which indicates magic injection that is below capacity. As the circuit depth increases, the behavior of $\overline{Y}$ approaches that of a Haar random $U_{A}$.

### v3

@@ C.2 A useful quantity and its approximations
where $P_{A}$ is a Pauli string supported on subsystem $A$. In the above definition, we also abuse notation by using $\mathbb{S}_{A}$ to represent an equal weight summation of all elements in $\mathbb{S}_{A}$: $\sum_{g\in\mathbb{S}_{A}}g$. The connection between the above quantity and $\overline{Y^{\rm lin}}$ will be explained in detail in the next subsection. Basically it naturally arises when computing ${\rm Tr}(PU_{A}\rho U_{A}^{\dagger})^{4}$, with $\rho$ represented as a projector onto the stabilizer subspace.
Eq. (14) can be represented using the diagrammatic method introduced in Sec. C.1 as follows:
Figure 8: (a) Leading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$. There are ${4!}^{2}=576$ diagrams in total. If $P_{A}\neq I_{A}$, then $\text{Tr}(P_{A})=\text{Tr}(P_{A}^{3})=0$, and there are 216 diagrams with non-zero contributions. (b) Subleading order and the corresponding diagrams for $\alpha_{P_{A}}(u_{1},u_{2},u_{3},u_{4})$.
@@
C.3 Derivation of the key results presented in the main text
In this subsection, we provide a detailed derivation of Eqs. (2), (LABEL:eq:UaUb) in the main text, as well as Eq. (9) in Sec. B. Since what we essentially need to compute is $1-\overline{Y^{\rm lin}}$, we will focus on this part below.
C.3.1 Eq. (2)
We start with Eq. (2). The key to our derivation is to represent the initial state as a projector onto the stabilizer subspace, and a decomposition of the stabilizer group of the full system in terms of the cosets of stabilizer subgroups of subsystems $A$ and $B$. Let $\mathbb{S}_{AB}$ denote the stabilizer group of the initial state. Then initial density matrix can be written as a projector
@@ C.3 Derivation of the key results presented in the main text
where $a_{i}\notin\mathbb{S}_{A}$ and $b_{i}\notin\mathbb{S}_{B}$ are the logical operators of subsystems $A$ and $B$, respectively. To ensure the proper commutation relation, the logical operators must be paired up into $(a_{i},b_{i})$ such that: (i) if $a_{i},a_{j}]=0$, then $b_{i},b_{j}]=0$; or (ii) if $\{a_{i},a_{j}\}=0$, then $\{b_{i},b_{j}\}=0$. While it is not hard to convince oneself that the above decomposition is correct for pure states with ${\rm dim}(\mathbb{S}_{AB})=N$, let us check that the dimensions work out. First of all, since the entanglement between $A$ and $B$ is $E$, there are precisely $E$ logical qubits encoded in both subsystems, with $4^{E}$ logical operators in each subsystem, and hence $4^{E}$ terms in the decomposition. Secondly, we can count the total number of group elements in $\mathbb{S}_{AB}$ from the right-hand side of the decomposition. We have $4^{E}\times 2^{{\rm dim}(\mathbb{S}_{A})}\times 2^{{\rm dim}(\mathbb{S}_{B})}=2^{|A|+|B|}=2^{N}$, as it should. We will give a concrete example of such a decomposition in Appendix D.
@@
Armed with the above representation of $\rho_{0}$, we can cast the average value $\overline{Y^{\rm lin}}$ as:
$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{A}}{\mathbb{S}_{AB}}{{U}_{A}}^{\dagger})}}$ (21)
@@
Consider first a fixed Pauli string $P$ and $k$, which corresponds to one term of the summation in $\overline{Y^{\rm lin}}$. We have the following result:
@@
[…]
@@
The qualitative features of Eq. (2), particularly the factor $2^{-|A|-E}$, can also be understood—since the unitary acts only on subsystem $A$, stabilizers supported entirely on $B$ remain unchanged under $U_{A}$ and contribute to $1-Y^{\rm lin}(\rho)$. A simple counting shows that the number of surviving stabilizers is precisely $2^{|B|-E}$, leading to the factor $2^{-|A|-E}$ in Eq. (2). The prefactor 4 in front of $2^{-|A|-E}$ is, in some sense, non-universal. To see this, let us define the generalized linear stabilizer entropy $Y_{\alpha}^{\rm lin}(\rho):=1-2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2\alpha}$. Eq. (2) thus corresponds to the specific case $\alpha=2$. Similarly, $1-Y_{\alpha}^{\rm lin}(\rho)$ is related to the $\alpha$-th stabilizer Renyi entropy: $M_{\alpha}(\rho)=\frac{1}{1-\alpha}{\rm log}(1-Y_{\alpha}^{\rm lin}(\rho))$. In the limit $\alpha\rightarrow\infty$, only the surviving stabilizers supported on $B$ contribute, yielding $\overline{Y_{\infty}^{\rm lin}}=1-2^{-|A|-E}$. For arbitrary $\alpha$, we have $1-\overline{Y_{\alpha}^{\rm lin}}=[1+c_{\alpha}(|A|,E)]2^{-|A|-E}$, with the one coming from the surviving stabilizers, and $c_{\alpha}(|A|,E)$ coming from contributions of all other Pauli strings, which in general depends on both $|A|$ and $E$  [^1] We thank Xhek Turkeshi for pointing this out to us.. Specifically, Eq. (2) implies that for $\alpha=2$, we have $c_{2}=3$.
C.3.2 Eq. (LABEL:eq:UaUb)
We provide detailed calculations for Eq. (LABEL:eq:UaUb) in the main text. Since the unitary acting on $\rho_{0}$ takes a factorized form $U=U_{A}\otimes U_{B}$, the average $\overline{Y^{\rm lin}}$ similarly factorizes into
$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}{{{\alpha}_{{{P}_{A}}}}({{a}_{{{k}_{1}}}},{{a}_{{{k}_{2}}}},{{a}_{{{k}_{3}}}},{{a}_{{{k}_{4}}}}){{\beta}_{{{P}_{B}}}}({{b}_{{{k}_{1}}}},{{b}_{{{k}_{2}}}},{{b}_{{{k}_{3}}}},{{b}_{{{k}_{4}}}})}},$ (24)
@@
We now analyze the order of magnitude of each term. Since $\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}\cdot(...)\leq\sum\limits_{{{k}_{1}},{{k}_{2}},{{k}_{3}},{{k}_{4}}}f_{i}$, where ()(...) represents any product of $\delta_{1},\delta_{2}$, or $\delta_{4}$ (as additional $\delta$-functions further reduce the number of unconstrained summations over $k_{1},\ldots,k_{4}$), we have
@@
After combining all these contributions and including the multiplicity factors coming from the number of Pauli strings in each case, we obtain the final expression in Eq. (LABEL:eq:UaUb) of the main text.
C.3.3 Eq. (9)
We now present the calculation details of Eq. (9) in Sec. B. Let us first recall the setup we use for this case. We consider a tripartite system $A$, $B$, and $C$. Since any tripartite stabilizer state can be transformed into a set of single-qubit states, Bell states, and GHZ states via local Clifford unitaries, we consider initial state consisting of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, ACAC, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem, shown in Fig. 5(a). Let us start by explicitly writing down a set of stabilizer generators for this state.
@@
where $l_{AB}^{k}$ and $l_{C}^{k}$ are logical operators on $AB$ and $C$ as before. In this case, since we are dealing with initial state with a factorized structure, it is easy to directly identify the logical operators. For example, $\{l_{AB}^{k}\}$ consists of Pauli operators acting on $b_{AC}+b_{BC}$ Bell pairs, as well as Pauli operators acting on $g$ GHZ states, hence $4^{b_{AC}+b_{BC}+g}$ operators in total. As before, to ensure elements across different cosets have the correct commutation relation as stabilizer group elements, we require (i) if $l_{AB}^{i},l_{AB}^{j}]=0$, then $l_{C}^{i},l_{C}^{j}]=0$; or (ii) if $\{l_{AB}^{i},l_{AB}^{j}\}=0$, then $\{l_{C}^{i},l_{C}^{j}\}=0$.
The average $\overline{Y^{\rm lin}}$ can be written as
$1-\overline{Y^{\rm lin}}=\frac{1}{{{2}^{5N}}}\sum\limits_{P\in P_{N}}{\overline{{{\text{Tr}}^{4}}(P{{U}_{AB}\otimes I_{C}}{\mathbb{S}_{ABC}}{U}_{AB}^{\dagger}\otimes I_{C})}}$ (34)
@@
The $\overline{Y^{\rm lin}}$ is then expressed in Eq. (61):
$2^{5|A|+5|B|+5|C|}\cdot 2^{-4|C|}\cdot 2^{-|C|+b_{AC}+b_{BC}+g}\cdot(1-\overline{Y^{\rm lin}})$ (61)
@@
Appendix D An example of the decomposition of the stabilizer group
@@
$\quad\quad\mathbb{S}_{A}\otimes\mathbb{S}_{B}\quad\quad$ $XXX\mathbb{S}_{A}\otimes XX\mathbb{S}_{B}$ $IIZ\mathbb{S}_{A}\otimes ZI\mathbb{S}_{B}$ $XXY\mathbb{S}_{A}\otimes YX\mathbb{S}_{B}$
@@
Appendix E Factorized unitary acting on three subregions
@@
$\overline{Y^{\rm lin}}=1-4\cdot 2^{-N}\big[1+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{AC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AB}}-2{{b}_{BC}}-2{g}}}+3\cdot{{2}^{-2{{b}_{AC}}-2{{b}_{BC}}-2{g}}}$ (66)
@@
where once again subleading corrections are of the order $O(2^{-|A|},2^{-|B|},2^{-|C|})$. Similarly to the bipartite result Eq. (LABEL:eq:UaUb), we find that the resulting state also achieves maximal amount of magic, provided that the subsystems are sufficiently entangled. Interestingly, this doesn’t require that the $A$, $B$, and $C$ are pairwise entangled. As Eq. (66) suggests, if A&BA\&B, B&CB\&C are highly entangled while A&CA\&C are not entangled (i.e. $b_{AB},b_{BC}$ is large, but $b_{AC}$ is zero), the average $\overline{Y^{\rm lin}}$ still approaches that of a Haar random state.
As a consistency check, we consider a special case of Eq. (66), namely, when only A&BA\&B are entangled, and $C$ is not entangled with $AB$. In this case, we can check that the average $1-\overline{Y^{\rm lin}}$ coincides with a product of Eq. (LABEL:eq:UaUb) in the main text for the bipartite case and a single Haar random state on $C$, which is consistent with our previous results.
@@
$1-\overline{Y^{\rm lin}}=\frac{1}{{2^{5N}}}\sum\limits_{P\in P_{N}}\sum\limits_{k_{1},k_{2},k_{3},k_{4}}\alpha_{P_{A}}(a_{k_{1}},a_{k_{2}},a_{k_{3}},a_{k_{4}})\beta_{P_{B}}(b_{k_{1}},b_{k_{2}},b_{k_{3}},b_{k_{4}})\gamma_{P_{C}}(c_{k_{1}},c_{k_{2}},c_{k_{3}},c_{k_{4}}).$ (68)
Now, depending on the operator contents of the Pauli string $P$ on subsystems $A$, $B$ and $C$ (identity or not), we use Eq. (16) and Eq. (17) to evaluate the above expression. We give an example here. Consider when $P_{A}\neq I_{A}$, $P_{B}\neq I_{B}$, and $P_{C}\neq I_{C}$, using Eq. (16) and Eq. (17), we expand Eq. (68) as follows: